# Limits on the Statistical Description of Charged de Sitter Black Holes

### v1

@@ 3.1 Heat Capacity at Fixed Charge
which agrees with (3.3). The heat capacity goes to zero for small (Planck mass) black holes, as it should. Concretely, in the semi-classical limit $\sqrt{G_{4}}/\ell_{4}\ll 1$ we find that $\tilde{C}_{b}=-1$ for Planck mass black holes with radius $r_{b}\sim\sqrt{G_{4}}$. Hence, general relativity breaks down in an ‘ordinary’ way for these black holes. We conclude that, expressed in terms of the redshifted energy $\tilde{M}$, in the Nariai limit there are still a large number of degrees of freedom available that can absorb heat and the statistical description does not break down.
@@
This expression reveals some interesting physics. First, we notice that at $r_{b}=\ell_{4}/2$ the heat capacity $\tilde{C}_{b}^{Q}$ diverges. This point corresponds to the intersection of the lukewarm line and the Nariai line at ${\cal Q}/\ell_{4}=1/4$. A divergent heat capacity is indicative of a phase transition [29] (however, see [30]). Indeed, the qualitative behavior of the black hole states change when crossing the lukewarm line. Black holes above the lukewarm line (in Fig. 1) have a temperature that is lower than that of the cosmological horizon. Below the lukewarm line, the heat capacity is negative and above the lukewarm line the heat capacity is positive, see Fig. 4. We will come back to the interpretation of this divergence in Sec. 3.2 and show how working in a different ensemble gives a finite and positive heat capacity.
@@
For temperatures smaller than this, we expect the thermodynamic description to break down.
@@ 3.2 Emission of Charged Particles
Up to now, we only studied the heat capacity of black holes at fixed charge. This is appropriate to study the thermodynamic behavior of black holes that only exchange (neutral) heat, but no charge. However, in general we expect the decay to be governed by (a combination of) Schwinger pair production and neutral Hawking radiation. [^5] It is particularly important to understand the possible decay channels of near-Nariai black holes through Schwinger pair production [9, 22]. When this is the case, we need a different ensemble to describe the thermodynamics of the black hole system.
Two cases where we already encountered this were the intersection of the lukewarm and Nariai line and the ultracold point. In the first case, we found that the heat capacity at fixed charge diverges. [^6] As we will show in more detail later, the heat capacity at fixed charge at the lukewarm line is finite in general. The heat capacity only diverges at the intersection of the lukewarm and Nariai line. In the second case, variations away from the ultracold point require variations in both the charge and mass to remain within the sharkfin.
@@ 3.2.1 Heat Capacity at Fixed Charge-to-Mass Ratio
As mentioned, the existence of the Schwinger effect necessitates the discussion of black hole thermodynamics in an ensemble where charge can vary. We will be interested in the heat capacity under Bousso-Hawking normalization, but with the constraint $\delta Q/\delta M=z$ when taking variations. [^7] We could also have expressed this constraint in terms of $\delta Q/\delta\tilde{M}$ which would introduce an additional redshift factor. Under this constraint, heat flow through both horizons equals
@@
Note that the combination $z\Phi_{b,c}$ appearing here does not have a tilde. For uncharged radiation ($z=0$) the heat flow through the black hole horizon always equals minus the heat flow through the cosmological horizon. Including charge, this is no longer true. We can for example consider the situation where there is positive (or vanishing) heat flow through both horizons, which occurs when
@@ 3.3 A Detailed Look at the Lukewarm Line
Because the Hawking temperature of the black hole and the cosmological horizon is the same, one would expect the lukewarm black holes to be in equilibrium when the spacetime is filled with Hawking radiation. This equilibrium state will be broken, however, if the Schwinger effect can happen, subsequently putting the system in an ensemble satisfying $\delta\mathcal{Q}/\delta\mathcal{M}=\mathcal{Z}$, where the $\mathcal{Z}$ is determined by the charged particle in question. [^8] The fixed charge-to-mass ratio constraint can be written in terms of $z=\delta Q/\delta M$ or $\mathcal{Z}=\delta{\cal Q}/\delta{\cal M}$. We’ll use them interchangeably.
In fact, the Schwinger effect always happen for lukewarm black holes for any charged particle spectrum, unlike cold black holes with a threshold value for the charge-to-mass ratio. The presence of a threshold charge-to-mass ratio for the Schwinger effect to take place has been used as a motivation for the Weak Gravity Conjecture [32, 33]. The absence of a threshold value can be seen from the near-horizon structure of the lukewarm black hole.
@@
where we made the change of coordinates $\rho=\frac{C}{4}x^{2}$ and defined the acceleration $\alpha=\frac{C}{2}$. The above metric describes the local ${\rm Rindler}_{2}\times S^{2}$ space. Similarly, the near-horizon geometry of the cosmological horizon can also be identified to be ${\rm Rindler}\times S^{2}$ with a different $S^{2}$ size.
With the coordinate expansion, the gauge field near the black hole horizon can be expressed as
@@
where $A_{t}^{(0)}$ is a constant piece. Since $F=\mathrm{d}A$ is proportional to the two-dimensional volume form $\sqrt{-g_{2}}\mathrm{d}^{2}x$, there is a constant electric field $E=\frac{Q}{4\pi r_{b}^{2}}$ in the near-horizon two-dimensional Rindler space. The same holds true at the cosmological horizon again with a weaker electric field strength.
@@
Notably, the above implies that, unlike extremal black holes whose Schwinger effect see a threshold value in the charge-to-mass ratio of the produced particle, lukewarm black holes emit charged radiation regardless of the particle spectrum. Depending on the charge-to-mass ratio of the particle, the Schwinger effect can either take the black hole above or below the lukewarm line, where Hawking radiation then takes action and drives the system back to the lukewarm solution.
@@
$\begin{cases}\delta\mathcal{M}=\frac{r_{b}^{2}(r_{b}^{2}-\ell_{4}^{2})-r_{c}^{2}(r_{c}^{2}-\ell_{4}^{2})+2(2r_{b}^{2}-\ell_{4}^{2})r_{b}(r_{c}-r_{b})}{2\ell_{4}^{2}(r_{c}-r_{b})^{2}}\delta r_{b}-\frac{r_{b}^{2}(r_{b}^{2}-\ell_{4}^{2})-r_{c}^{2}(r_{c}^{2}-\ell_{4}^{2})+2(2r_{c}^{2}-\ell_{4}^{2})r_{c}(r_{c}-r_{b})}{2\ell_{4}^{2}(r_{c}-r_{b})^{2}}\delta r_{c}\\ \delta(\mathcal{Q}^{2})=\frac{1}{(r_{c}^{-1}-r_{b}^{-1})^{2}\ell_{4}^{2}}\left\{\left[(-3r_{b}^{2}+\ell_{4}^{2})(r_{c}^{-1}-r_{b}^{-1})-\frac{r_{c}(r_{c}^{2}-\ell_{4}^{2})-r_{b}(r_{b}^{2}-\ell_{4}^{2})}{r_{b}^{2}}\right]\delta r_{b}\right.\\ \left.-\left[(-3r_{c}^{2}+\ell_{4}^{2})(r_{c}^{-1}-r_{b}^{-1})-\frac{r_{c}(r_{c}^{2}-\ell_{4}^{2})-r_{b}(r_{b}^{2}-\ell_{4}^{2})}{r_{c}^{2}}\right]\delta r_{c}\right\}\end{cases}.$ (3.21)
@@
While the quantities evaluated at $r=r_{b}$ can be simplified after restricting to the lukewarm line, the heat capacity in Bousso-Hawking normalization generally requires numerical evaluation due to the observer location $r_{\cal O}$ being the root of a quartic equation, which has a lengthy expression in general. The behavior of $\tilde{C}^{z}_{b}$ with different choices of charge-to-mass ratio is shown in Fig. 5.
Figure 5: Heat capacity of lukewarm black holes in various ensembles specified by $\mathcal{Z}$. From top to bottom, the three curves correspond to $\mathcal{Z}=2.2,2,1.8$ respectively. When $\mathcal{Z}>2$, $\tilde{C}^{z}_{b}\rightarrow\infty$ at the lukewarm-Nariai intersection point where as for $\mathcal{Z}<2$, $\tilde{C}^{z}_{b}\rightarrow$-$\infty$. For the later case, $\tilde{C}^{z}_{b}$ has a zero at the point $1-\frac{\mathcal{Z}r_{c}}{\ell_{4}}=0$ which corresponds to $\delta S_{b}=0$.
From (3.10), we saw that $\delta S_{b}=0$ when $z=\Phi_{b}^{-1}$ Along the lukewarm line, which corresponds to $z=2\sqrt{\pi G_{4}}\ell_{4}r_{c}^{-1}$. The consequence is that at this point the heat capacity also vanishes. However, this result does not directly indicate a breakdown of thermodynamics since we only allow decay channels that change the black hole charge-to-mass ratio in a constrained manner. In the canonical ensemble (fixed charge), a vanishing heat capacity reflects that the change in mass becomes independent of the temperature. Schwinger pair production on the other hand, does not depend directly on temperature. A vanishing heat capacity at fixed $z$ therefore does not reflect a similar breakdown of thermodynamics.
@@
One intriguing aspect of working at fixed $z$ is that it allows one to tune the approach to the intersection of the lukewarm and Nariai line. As already mentioned, working at fixed charge this point has a diverging heat capacity suggestive of a phase transition when crossing the lukewarm line. By tuning $z$ to obtain a finite value of the heat capacity, it is perhaps possible to better characterize the properties of this transition.
@@ 4 Near-Extremal Expansion
Before we conclude, in this section we will give a complete comparison of the different thermodynamic quantities using the $\gamma=1$ and Bousso-Hawking normalization. We will focus on the different near-extremal limits ((charged) Nariai, cold) in an ensemble of fixed charge and tabulate the behavior of various thermodynamic quantities in Tab. 3. The ultracold case will be treated separately.
@@
The thermodynamic quantities coming from the expansions above are presented in Tab. 3, in agreement with our previous results. We emphasize that $\epsilon$ can be expressed in terms of any (dimensionless) near-extremal thermodynamic quantity.
@@
where again rescaled $r_{b1}$ without any loss of generality and consistent with $\epsilon\geq 0$ and dimensionless. On top of requiring $f^{\prime}(r_{\mathcal{O}})=0+\mathcal{O}(\epsilon^{4})$ and $\delta\ell_{4}=\mathcal{O}(\epsilon^{4})$ to fix the coefficients of $r_{\mathcal{O}}$ and $r_{c}$, we impose $\delta Q=\mathcal{O}(\epsilon^{4})$ in order to fix the coefficients of $r_{a}$ and enforce the canonical ensemble. The explicit expressions for these coefficients are straightforward to obtain, but large and not very insightful. We will therefore not present them here. [^9] We are happy to provide them upon request. After using the expansions above as input we obtain the thermodynamic quantities as presented in Tab. 3.
@@
The results are displayed in Tab. 3, but we denote a few of the more lengthy quantities separately here.
@@
From the above results, we can deduce where the thermodynamic description breaks down due to the (absolute value of the) heat capacity becoming smaller than one. Whenever this happens, one might expect large (quantum) corrections to modify the behavior of the partition function, as is the case for log-$T$ corrections.
@@
For Nariai black holes, we found that sufficiently far away from the ultracold point the (absolute value of the) heat capacity remains large when using Bousso-Hawking normalization. This signals no breakdown of the thermodynamic description, therefore questioning the presence of large (quantum) corrections. This should be contrasted with $\gamma=1$ normalization where the heat capacity is zero along the entire Nariai branch. It would therefore be of great interest to compute the one-loop determinant of near-Nariai black holes using Bousso-Hawking normalization to see if this intuition is correct and whether the normalization of the Killing vector plays an important role. Existing results on one-loop determinants for de Sitter black holes include [34, 35, 36, 37, 28, 38, 39] and some of these references did report large quantum corrections in the Nariai limit (see e.g. [36, 37, 28] for recent results). These results, however, seem to depend on a complexification of the geometry describing the Milne patch of the near-horizon de Sitter region, instead of the static patch that we focus on. Therefore, these one-loop results apply in a different regime. Nonetheless, it would be interesting to better understand the relation between these computations and our thermodynamic approach.
Lastly, for ultracold black holes the value of the heat capacity is again equal for both choices of normalization so we do expect a breakdown of the thermodynamic decscription and, most likely, large corrections. Explicitly computing these corrections in the ultracold limit, however, has remained a challenge.

### v2

@@ 3.1 Heat Capacity at Fixed Charge
which agrees with (3.3). The heat capacity goes to zero for small (Planck mass) black holes, as it should. Concretely, in the semi-classical limit $\sqrt{G_{4}}/\ell_{4}\ll 1$ we find that $\tilde{C}_{b}=-1$ for Planck mass black holes with radius $r_{b}\sim\sqrt{G_{4}}$. Hence, general relativity breaks down in the standard expected way for these black holes. We conclude that, expressed in terms of the redshifted energy $\tilde{M}$, in the Nariai limit there are still a large number of degrees of freedom available that can absorb heat and the statistical description does not break down.
@@
This expression reveals some interesting physics. First, we notice that at $r_{b}=\ell_{4}/2$ the heat capacity $\tilde{C}_{b}^{Q}$ diverges. This point corresponds to the intersection of the lukewarm line and the Nariai line at ${\cal Q}/\ell_{4}=1/4$. A divergent heat capacity is indicative of a phase transition [32] (however, see [33]). As will be discussed later, this intersection is associated with a continuous phase transition. Indeed, the qualitative behavior of the black hole states change when crossing the lukewarm line. Black holes above the lukewarm line (in Fig. 1) have a temperature that is lower than that of the cosmological horizon. Below the lukewarm line, the heat capacity is negative and above the lukewarm line the heat capacity is positive, see Fig. 4666More generally one can show that at fixed charge below the lukewarm-Nariai intersection, a maximum black hole temperature point exists in between the lukewarm and the Nariai state. This is sometimes referred to as the Davies point [34]. Consequently, this corresponds to a transition from positive to negative heat capacity. These points form a line that intersects with the lukewarm and Nariai line exactly at the lukewarm-Nariai intersection.. We will come back to the interpretation of this divergence in Sec. 3.2 and show how working in a different ensemble gives a finite and positive heat capacity.
@@
For temperatures smaller than this, we expect the thermodynamic description to break down. As the dS2 section becomes infinitely large in the ultracold limit, this is probably connected to a strong coupling divergence in the IR of the reduced two-dimensional dilaton gravity description [21].
@@ 3.2 Emission of Charged Particles
Up to now, we only studied the heat capacity of black holes at fixed charge. This is appropriate to study the thermodynamic behavior of black holes that only exchange (neutral) heat, but no charge. However, in general we expect the decay to be governed by (a combination of) Schwinger pair production and neutral Hawking radiation [^7] It is particularly important to understand the possible decay channels of near-Nariai black holes through Schwinger pair production [9, 22].. When this is the case, we need a different ensemble to describe the thermodynamics of the black hole system.
We already encountered two cases where this might be relevant: the intersection of the lukewarm and Nariai line and the ultracold point. In the first case, we found that the heat capacity at fixed charge diverges [^8] As we will show in more detail later, the heat capacity at fixed charge at the lukewarm line is finite in general. The heat capacity only diverges at the intersection of the lukewarm and Nariai line.. In the second case, variations away from the ultracold point require variations in both the charge and mass to remain within the sharkfin.
@@ 3.2.1 Heat Capacity at Fixed Charge-to-Mass Ratio
As mentioned, the existence of the Schwinger effect necessitates the discussion of black hole thermodynamics in an ensemble where charge can vary. We will be interested in the heat capacity using Bousso-Hawking normalization, but with the constraint $\delta Q/\delta M=z$ when considering variations. [^9] We could also have expressed this constraint in terms of $\delta Q/\delta\tilde{M}$ which would introduce an additional redshift factor. Under this constraint, the heat flow through both horizons equals
@@
Note that the combination $z\Phi_{b,c}$ appearing here does not have a tilde. For neutral emissions ($z=0$) the heat flow through the black hole horizon always equals minus the heat flow through the cosmological horizon. Including charge, this is no longer true. We can for example consider the situation where there is positive (or vanishing) heat flow through both horizons, which occurs when
@@ 3.3 A Detailed Look at the Lukewarm Line
Because the Hawking temperature of the black hole and the cosmological horizon is the same, one would expect the lukewarm black holes to be in equilibrium when the spacetime is filled with Hawking radiation. This equilibrium will be broken by the emission of charged particles through the Schwinger effect, subsequently putting the system in an ensemble satisfying $\delta\mathcal{Q}/\delta\mathcal{M}=\mathcal{Z}$, where the $\mathcal{Z}$ is determined by the charged particle in question. [^10] The fixed charge-to-mass ratio constraint can be written in terms of $z=\delta Q/\delta M$ or $\mathcal{Z}=\delta{\cal Q}/\delta{\cal M}$. We’ll use them interchangeably.
In fact, the Schwinger effect always occurs for lukewarm black holes for any charged particle spectrum, unlike cold black holes where there exists a threshold value for the charge-to-mass ratio. The presence of a threshold charge-to-mass ratio has been used as a motivation for the Weak Gravity Conjecture [36, 37]. The absence of a threshold value can be seen from the near-horizon structure of the lukewarm black hole.
@@
where we made the change of coordinates $\rho=\frac{C}{4}x^{2}$ and defined the acceleration $\alpha=\frac{C}{2}$. The above metric describes the local ${\rm Rindler}_{2}\times S^{2}$ space. Similarly, the near-horizon geometry of the cosmological horizon can also be identified as ${\rm Rindler}\times S^{2}$, but with a different $S^{2}$ size.
Using the near-horizon expansion, the gauge field near the black hole horizon can be expressed as
@@
where $A_{t}^{(0)}$ is a constant piece. Since $F=\mathrm{d}A$ is proportional to the two-dimensional volume form $\sqrt{-g_{2}}\mathrm{d}^{2}x$, there is a constant electric field $E=\frac{Q}{4\pi r_{b}^{2}}$ in the near-horizon two-dimensional Rindler space. The same holds true at the cosmological horizon, just with a smaller electric field strength.
@@
Notably, the above implies that, unlike extremal black holes whose Schwinger effect feature a threshold value in the charge-to-mass ratio of the produced particle, lukewarm black holes emit charged radiation regardless of the particle spectrum. Depending on the charge-to-mass ratio of the particle, the Schwinger effect can either take the black hole above or below the lukewarm line, where neutral Hawking radiation can then drive the system back to the lukewarm solution.
@@
$\begin{cases}\delta\mathcal{M}=\frac{r_{b}^{2}(r_{b}^{2}-\ell_{4}^{2})-r_{c}^{2}(r_{c}^{2}-\ell_{4}^{2})+2(2r_{b}^{2}-\ell_{4}^{2})r_{b}(r_{c}-r_{b})}{2\ell_{4}^{2}(r_{c}-r_{b})^{2}}\delta r_{b}-\frac{r_{b}^{2}(r_{b}^{2}-\ell_{4}^{2})-r_{c}^{2}(r_{c}^{2}-\ell_{4}^{2})+2(2r_{c}^{2}-\ell_{4}^{2})r_{c}(r_{c}-r_{b})}{2\ell_{4}^{2}(r_{c}-r_{b})^{2}}\delta r_{c}\\ \delta(\mathcal{Q}^{2})=\frac{1}{(r_{c}^{-1}-r_{b}^{-1})^{2}\ell_{4}^{2}}\left\{\left[(-3r_{b}^{2}+\ell_{4}^{2})(r_{c}^{-1}-r_{b}^{-1})-\frac{r_{c}(r_{c}^{2}-\ell_{4}^{2})-r_{b}(r_{b}^{2}-\ell_{4}^{2})}{r_{b}^{2}}\right]\delta r_{b}\right.\\ \left.\qquad\qquad-\left[(-3r_{c}^{2}+\ell_{4}^{2})(r_{c}^{-1}-r_{b}^{-1})-\frac{r_{c}(r_{c}^{2}-\ell_{4}^{2})-r_{b}(r_{b}^{2}-\ell_{4}^{2})}{r_{c}^{2}}\right]\delta r_{c}\right\}\end{cases}.$ (3.21)
@@
While the quantities evaluated at $r=r_{b}$ can be simplified after restricting to the lukewarm line, the heat capacity in Bousso-Hawking normalization generally requires numerical evaluation due to the observer location $r_{\cal O}$ being the root of a quartic equation, which in general is a complicated lengthy expression. The behavior of $\tilde{C}^{z}_{b}$ with different choices of charge-to-mass ratio is shown in Fig. 5.
Figure 5: Heat capacity of lukewarm black holes in various ensembles specified by $\mathcal{Z}$. From top to bottom, the three curves correspond to $\mathcal{Z}=2.2,2,1.8$ respectively. When $\mathcal{Z}>2$, $\tilde{C}^{z}_{b}\rightarrow\infty$ at the lukewarm-Nariai intersection point whereas for $\mathcal{Z}<2$, $\tilde{C}^{z}_{b}\rightarrow$-$\infty$. For the latter case, $\tilde{C}^{z}_{b}$ has a zero at the point $1-\frac{\mathcal{Z}r_{c}}{\ell_{4}}=0$ which corresponds to $\delta S_{b}=0$.
From (3.10), we saw that $\delta S_{b}=0$ when $z=\Phi_{b}^{-1}$ along the lukewarm line, which corresponds to $z=2\sqrt{\pi G_{4}}\ell_{4}r_{c}^{-1}$. As a consequence the heat capacity also vanishes. However, this result does not directly indicate a breakdown of thermodynamics since we only allow decay channels that change the black hole charge-to-mass ratio in a constrained manner. In the canonical ensemble (fixed charge), a vanishing heat capacity reflects that the change in mass becomes independent of the temperature. Schwinger pair production on the other hand, does not depend directly on temperature. A vanishing heat capacity at fixed $z$ therefore does not reflect a similar breakdown of thermodynamics.
@@
One intriguing aspect of working at fixed $z$ is that it allows one to tune the approach to the intersection of the lukewarm and Nariai line. As already mentioned, working at fixed charge this point has a diverging heat capacity suggestive of a continuous phase transition. To see this, we set ${\cal Z}=1$ to define an ensemble of lukewarm black holes. This allows us to approach the lukewarm-Nariai intersection while staying on the lukewarm line and track the change in energy and temperature by evaluating the heat capacity. For continuous phase transitions, the heat capacity has a power law divergence, the power of which defines a critical exponent that can be used to distinguish universality classes. The temperature of lukewarm black holes near the intersection is expanded as
$\tilde{T}_{b}(r_{b})=\tilde{T}_{\text{crit}}\left[1-\left(\frac{r_{b}}{\ell_{4}}-\frac{1}{2}\right)^{2}+\cdots\right],$ (3.27)
and the entropy is
$S_{b}(r_{b})=\frac{\pi\ell_{4}^{2}}{4G_{4}}+\frac{\pi\ell_{4}^{2}}{G_{4}}\left(\frac{r_{b}}{\ell_{4}}-\frac{1}{2}\right)+\cdots.$ (3.28)
From these two expansions, we see that the heat capacity diverges as
$C_{\text{LW}}(r_{b})=\tilde{T}_{b}\left.\frac{\delta S_{b}}{\delta\tilde{T}_{b}}\right|_{\text{LW}}\propto\left(\frac{r_{b}}{\ell_{4}}-\frac{1}{2}\right)^{-1}~,$ (3.29)
such that we find
$C_{\text{LW}}\propto\left(\frac{\tilde{T}_{\text{crit}}-\tilde{T}_{b}}{\tilde{T}_{\text{crit}}}\right)^{-\frac{1}{2}},$ (3.30)
indicating a critical exponent of $\alpha=\frac{1}{2}$. This exponent is associated to the divergence of the macroscopic response function $\tilde{C}_{b}$ and provides part of the information about the universality class of the lukewarm-Nariai critical point. The full information requires also microscopics of the black holes near criticality, for instance, the order parameter, correlation function and correlation length of gravitational fluctuations.
@@ 4 Near-Extremal Expansion
Before we conclude, in this section we will give a complete comparison of the different thermodynamic quantities using the $\gamma=1$ and Bousso-Hawking normalization. We will focus on the different near-extremal limits ((charged) Nariai, cold) in an ensemble of fixed charge and tabulate the behavior of various thermodynamic quantities in Table 3. The ultracold case will be treated separately.
@@
The thermodynamic quantities coming from the expansions above are presented in Table 3, in agreement with our previous results. We emphasize that $\epsilon$ can be expressed in terms of any (dimensionless) near-extremal thermodynamic quantity.
@@
where again rescaled $r_{b1}$ without any loss of generality and consistent with $\epsilon\geq 0$ and dimensionless. On top of requiring $f^{\prime}(r_{\mathcal{O}})=0+\mathcal{O}(\epsilon^{4})$ and $\delta\ell_{4}=\mathcal{O}(\epsilon^{4})$ to fix the coefficients of $r_{\mathcal{O}}$ and $r_{c}$, we impose $\delta Q=\mathcal{O}(\epsilon^{4})$ in order to fix the coefficients of $r_{a}$ and enforce the canonical ensemble. The explicit expressions for these coefficients are straightforward to obtain, but complicated and not very insightful. We will therefore not present them here [^11] We are happy to provide them upon request.. After using the expansions above as input we obtain the thermodynamic quantities as presented in Table 3.
@@
The results are displayed in Table 3, but we denote a few of the more lengthy quantities separately here.
@@
From the above results, we can deduce where the thermodynamic description breaks down due to the (absolute value of the) heat capacity becoming smaller than one. Whenever this happens, one might expect large (quantum) corrections to modify the behavior of the partition function, for instance through log-$T$ corrections.
@@
For Nariai black holes, we found that sufficiently far away from the ultracold point the (absolute value of the) heat capacity remains large when using Bousso-Hawking normalization. This signals a consistent thermodynamic description and therefore the absence of large (quantum) corrections. This should be contrasted with $\gamma=1$ normalization where the heat capacity is zero along the entire Nariai branch. It would clearly be of great interest to compute the one-loop determinant of near-Nariai black holes using Bousso-Hawking normalization to see if this intuition is correct and whether the choice for the normalization of the Killing vector is important. Existing results on one-loop determinants for de Sitter black holes include [38, 39, 40, 41, 31, 42, 43] and some of these references did report large quantum corrections in the Nariai limit (see e.g. [40, 41, 31] for recent results). These results, however, seem to depend on a complexification of the geometry describing the Milne patch of the near-horizon de Sitter region, instead of the static patch that we focus on. Therefore, these one-loop results apply in a different regime. Nonetheless, it would be interesting to better understand the relation between these computations and our thermodynamic approach.
Lastly, for ultracold black holes the value of the heat capacity is, again, the same for both choices of normalization. As a consequence we do expect a breakdown of the thermodynamic decscription and, most likely, large corrections. Explicitly computing these corrections in the ultracold limit, however, has remained a challenge.